# Client-Centered Federated Learning for Heterogeneous EHRs: Use Fewer Participants to Achieve the Same Performance

### v4

@@ Abstract
The increasing volume of electronic health records (EHRs) presents the opportunity to improve the accuracy and robustness of models in clinical prediction tasks. Unlike traditional centralized approaches, federated learning enables training on data from multiple institutions while preserving patient privacy and complying with regulatory constraints. In practice, healthcare institutions (i.e., hosts) often need to build predictive models tailored to their specific needs using federated learning. In this scenario, two key challenges arise: (1) ensuring compatibility across heterogeneous EHR systems, and (2) managing federated learning costs within budget constraints. To address these challenges, we propose EHRFL, a federated learning framework designed for building a cost-effective, host-specific predictive model using patient EHR data. EHRFL consists of two components: (1) text-based EHR modeling, which facilitates cross-institution compatibility without costly data standardization, and (2) a participant selection strategy based on averaged patient embedding similarity to reduce the number of participants without degrading performance. Experiments on multiple open-source EHR datasets demonstrate the effectiveness of both components. We believe our framework offers a practical solution for enabling healthcare institutions to build institution-specific predictive models under budgetary constraints.
@@ 1 Introduction
In such a scenario, the host faces two key challenges: (1) ensuring compatibility of EHR modeling across institutions with differing EHR systems, and (2) managing overall federated learning costs to stay within budgetary constraints. To address these challenges, we propose EHRFL, a framework designed for cost-effective federated learning to build targeted predictive models on EHR data for the host institution. Our framework comprises the following two components.
@@
Second, we propose a method for selecting suitable federated learning participants based on averaged patient embedding similarity, to reduce the number of participants without compromising host model performance. As the number of federated learning participants increases, the host incurs higher costs (e.g., data usage fees, computational costs) [14]. Thus, it is essential to effectively reduce the number of participating subjects without sacrificing the host’s model performance. Importantly, among the candidate subjects, certain subjects may not contribute positively or may even degrade the model performance for the host. Identifying and excluding such subjects is crucial for cost-effective federated learning. To address this, our framework introduces a method for selecting participating subjects based on similarity between the host’s and each subject’s averaged patient embeddings. Our approach aims to reduce the number of participating subjects while preserving host model performance, ultimately lowering the overall federated learning costs for the host.
In our experiments, we evaluate the two components of our framework. First, we show that text-based EHR federated learning enables collaborative training between institutions with heterogeneous EHR systems, achieving performance gains while maintaining cost-effectiveness. Second, we observe a correlation between performance gains for the host and average patient embedding similarity, supporting its use as a criterion for excluding subjects unlikely to contribute positively to the host model. Third, we validate that averaged patient embedding similarity-based selection effectively reduces the number of participants without negatively impacting host model performance. Overall, we believe our proposed EHRFL framework significantly enhances the practical applicability of federated learning in clinical settings, enabling healthcare institutions to build cost-effective, institution-specific predictive models within budget constraints.
@@ 2 Overall Framework
We introduce the two key components of our EHRFL framework. The two components are (1) text-based EHR federated learning (Figure 1) and (2) participating client selection using averaged patient embeddings (Figure 2).
@@
Figure 2: Selection of participating subjects in federated learning based on averaged patient embedding similarity with the host. To ensure privacy, neither raw data nor individual patient embeddings are shared; only averaged embeddings are shared. Subjects with low similarity scores relative to the host are excluded from the federated learning process.
@@ 2.1.2 Text-based EHR Linearization
$r_{i}=(e_{i}\oplus n_{i1}\oplus v_{i1}^{\prime}\oplus n_{i2}\oplus v_{i2}^{\prime}\cdots)$
@@ 2.2 Participating Subject Selection using Averaged Patient Embeddings
In this section, we present our method for selecting suitable participating subjects for the host using averaged patient embeddings. Motivated by prior research [15, 16, 17, 18], we hypothesize that subjects whose data distributions differ significantly from that of the host are more likely to have no positive impact or may even degrade the host model’s performance. To identify suitable subjects, we compute the similarity between the averaged patient embeddings extracted from the text-based EHR data of each candidate subject and that of the host.
@@ 2.2.1 Subject Selection Process
The selection process involves the following steps. First, the host trains a model on its own data to build a common model for extracting patient embeddings. This model is trained using the text-based EHR modeling method described in Section 2.1. This is to ensure that embeddings are generated within a shared embedding space, despite that host and subjects may have distinct EHR systems. Second, the host shares the trained model weights to each candidate subject. Third, the host and each candidate subject use the model to extract patient embeddings from their respective datasets. The resulting embeddings for the host are denoted as ($l^{h}_{1}$, $l^{h}_{2}$, …, $l^{h}_{n}$), while the resulting embeddings for a single candidate subject are denoted as ($l^{s}_{1}$, $l^{s}_{2}$, …, $l^{s}_{m}$), where n and m represent the number of patients for the host and subject, respectively (note that $l$ is consistent with the notation in Section 2.1.3). Fourth, the host and each candidate subject each generate an averaged patient embedding: the host generates $\bar{l^{h}}=\frac{1}{n}\sum_{i=1}^{n}l^{h}_{i}$, and each candidate subject generates $\bar{l^{s}}=\frac{1}{m}\sum_{i=1}^{m}l^{s}_{i}$. This averaged embedding results in a single embedding vector of dimensionality $d$, where $d$ is the dimensionality of each patient embedding $l$. Fifth, each candidate subject shares their averaged patient embedding, $\bar{l^{s}}$, to the host. Sixth, for each candidate subject, the host computes the similarity between $\bar{l^{h}}$ and $\bar{l^{s}}$. Similarity can be measured using metrics such as cosine similarity, euclidean distance, and KL divergence. Seventh, the host excludes candidate subjects with low similarity scores, proceeding federated learning with subjects of high similarity scores that are not excluded.
@@
A potential concern with this approach is that averaging patient embeddings across clients may result in information loss, which may lead to suboptimal selection of suitable clients. Specifically, given that each client may have many patients, measuring similarity using averaged patient embeddings may lack the necessary granularity. To address this concern, we additionally evaluate subject selection based on patient-wise similarities using non-averaged patient-level embeddings (see Appendix F). As detailed in Appendix F, our experiment results show that the averaged patient embedding method consistently leads to better host performance compared to selecting subjects using non-averaged embedding similarities. Moreover, sharing patient-level embeddings across institutions may raise privacy concerns (see Section 2.2.2 for details), making the average embedding-based approach a reasonable method from both performance and privacy perspectives.
2.2.2 Privacy Analysis
A potential concern when sharing averaged patient embeddings with the host is the risk of privacy leakage. Our method, however, is closely aligned with the principles of differential privacy [19, 20], which is widely considered the gold standard for privacy protection. Specifically, differential privacy permits the sharing of group-level (e.g., hospital-level) statistical information, provided that individual identities cannot be re-identified. In accordance with this principle, our approach averages embeddings across a large number of patients, significantly reducing data granularity and making it nearly impossible to infer individual-level information.
For scenarios requiring formal privacy guarantees, standard differential privacy techniques, such as adding calibrated noise to patient embeddings, can be applied to our method. Yet, even in this case, the appropriate amount of noise required to achieve complete privacy still remains debatable. We leave the integration of formal privacy mechanisms to future work, as our research focuses on exploring initial strategies for selecting suitable clients for the host.
@@ 2.3 Overall Cost Savings for the Host
Cost savings arise from avoiding data usage and computational costs for unselected clients. Specifically, as shown in Table 1, when only $K$ participants are involved instead of all $N$, both data usage fees and training costs are reduced proportionally by $(N-K)$. Among the participating client selection costs, the dominant factors are the data usage fee $X$ and the training cost $C_{\text{train}}$, as the other components—$C_{\text{model}}$, $C_{\text{extract}}$, $C_{\text{average}}$, $C_{\text{latent}}$, and $C_{\text{sim}}$—are associated with lightweight model inference or computation and are typically negligible in practice.
As shown in the total net savings equation, when $K\ll N$, cost savings are substantial since $X$ and $C_{\text{train}}$ scale with $(N-K)$. Even in the minimal case where $K=N-1$, the host still saves approximately $X+C_{\text{train}}$, indicating that our method provides cost benefits even under conservative selection scenarios.
@@
$N$ Total number of candidate clients in FL (including the host)
@@
Extracting latents $N\cdot C_{\text{extract}}$
Averaging latents $N\cdot C_{\text{average}}$
Transmitting latents to the host $N\cdot C_{\text{latent}}$
@@
Overall Costs $E\cdot C_{\text{train}}+N\cdot(C_{\text{model}}+C_{\text{extract}}+C_{\text{average}}+C_{\text{latent}}+C_{\text{sim}})$
@@
Total Net Savings $(N-K)\cdot X+((N-K)\cdot R\cdot L-E)\cdot C_{\text{train}}$ $+(2\cdot(N-K)\cdot R-N)\cdot C_{\text{model}}$ $-N\cdot(C_{\text{extract}}+C_{\text{average}}+C_{\text{latent}}+C_{\text{sim}})$
@@ 3.1 Datasets
In summary, we construct a total of five clients, representing four distinct EHR database systems, covering three different geographical regions. Detailed demographic information and label distributions for each client are provided in Appendix B.
@@ 3.2.1 Cohort
3.2.2 Prediction Tasks
We evaluate model performance on 12 clinical prediction tasks, following the tasks in [2]. A diverse set of tasks is included to demonstrate the effectiveness of our framework in building host models across various clinical scenarios. The default prediction window is 24 hours, unless otherwise specified (see Appendix A for detailed task descriptions). Each model is trained using a multi-task learning, with 12 distinct prediction heads each corresponding to the individual tasks.
@@ 3.2.4 Training Procedure
Since the objective of our framework is to develop a model tailored specifically for the host, the final host model selected is the one that achieves the highest validation performance for the host across all communication rounds. The experiments were conducted using four NVIDIA RTX 3090 GPUs with CUDA version 11.4, with each host model requiring approximately 1–2 days for training.
@@ 4 Results
Table 2: Comparison of the host performance (average macro AUROC of the 12 tasks) trained solely on local data (i.e., Single) versus trained on federated learning using text-based EHR modeling with a fixed number of clients. The standard deviation results are listed in Appendix C. Each client column represents the performance when the client is the host. Average and Best represent the average and best host performance across all subject combinations for the fixed number of clients. Bold indicates federated learning performance higher than Single results.
@@ 4.1 Federated Learning with Text-based EHR Modeling
In this section, we evaluate the effectiveness of federated learning using text-based EHR modeling. Table 2 shows a comparison of the host model performance when trained solely on its own data (i.e., Single) versus when trained using federated learning with text-based EHR modeling. The standard deviation results for Table 2 are listed in Appendix C. The performance of each host show that federated learning with text-based EHR modeling generally leads to improved host performance compared to training solely on local data (see bolded values). These results indicate that text-based EHR federated learning enables leveraging data from multiple clients to improve the host’s model performance, despite heterogeneity in EHR systems. This trend is consistent across all of the four federated learning algorithms evaluated. Importantly, text-based EHR modeling does not require extensive preprocessing or expert-driven domain-specific mappings, enabling low-cost federated learning when building a model for the host using heterogeneous EHR systems.

### v5

# Federated Learning for Heterogeneous Electronic Health Record Systems with Cost Effective Participant Selection

@@ Abstract
The increasing volume of electronic health records (EHRs) presents the opportunity to improve the accuracy and robustness of models in clinical prediction tasks. Unlike traditional centralized approaches, federated learning enables training on data from multiple institutions while preserving patient privacy and complying with regulatory constraints. In practice, healthcare institutions (i.e., hosts) often need to build predictive models tailored to their specific needs (e.g., creatinine-level prediction, N-day readmission prediction) using federated learning. When building a federated learning model for a single healthcare institution, two key challenges arise: (1) ensuring compatibility across heterogeneous EHR systems, and (2) managing federated learning costs within budget constraints. Specifically, heterogeneity in EHR systems across institutions hinders compatible modeling, while the computational costs of federated learning can exceed practical budget limits for healthcare institutions. To address these challenges, we propose EHRFL, a federated learning framework designed for building a cost-effective, host-specific predictive model using patient EHR data. EHRFL consists of two components: (1) text-based EHR modeling, which facilitates cross-institution compatibility without costly data standardization, and (2) a participant selection strategy based on averaged patient embedding similarity to reduce the number of participants without degrading performance. Our participant selection strategy sharing averaged patient embeddings is differentially private, ensuring patient privacy. Experiments on multiple open-source EHR datasets demonstrate the effectiveness of both components. With our framework, healthcare institutions can build institution-specific predictive models under budgetary constraints with reduced costs and time.
@@ 1 Introduction
In such a scenario, the host faces two key challenges: (1) ensuring compatibility of EHR modeling across institutions with differing EHR systems [2], and (2) managing overall federated learning costs to stay within budgetary constraints [10, 11]. Addressing both EHR heterogeneity and modeling cost constraints is critical for practical deployment of federated learning in healthcare [2, 10, 11]. To address these challenges, we propose EHRFL, a framework designed for cost-effective federated learning to build targeted predictive models on EHR data for the host institution. Our framework comprises the following two components.
@@
Second, we propose a method for selecting suitable federated learning participants based on averaged patient embedding similarity, to reduce the number of participants without compromising host model performance. As the number of federated learning participants increases, the host incurs higher costs (e.g., data usage fees, computational costs) [16]. Thus, it is essential to effectively reduce the number of participating subjects without sacrificing the host’s model performance. Importantly, among the candidate subjects, certain subjects may not contribute positively or may even degrade the model performance for the host. Identifying and excluding such subjects is crucial for cost-effective federated learning. To address this, our framework introduces a method for selecting participating subjects based on similarity between the host’s and each subject’s averaged patient embeddings. To ensure patient privacy during the sharing averaged patient embeddings, we incorporate differential privacy [17] when constructing averaged patient embeddings used for selecting participants. Our approach aims to reduce the number of participating subjects while preserving host model performance, ultimately lowering the overall federated learning costs for the host.
In our experiments, we evaluate the two components of our framework. First, we show that text-based EHR federated learning enables collaborative training between institutions with heterogeneous EHR systems, achieving performance gains while maintaining cost-effectiveness. Second, we observe a correlation between performance gains for the host and average patient embedding similarity, supporting its use as a criterion for excluding subjects unlikely to contribute positively to the host model. Third, we validate that averaged patient embedding similarity-based selection effectively reduces the number of participants without negatively impacting host model performance. Our experimental results show that our framework enables the construction of an institution-specific model within budget constraints, with minimal cost and data-processing time. As discussed in the Discussions section, our work is limited by the use of ICU data only, due to resource constraints. In future work, we plan to extend our framework to other clinical settings (e.g., outpatient data, emergency department data). Overall, we believe our proposed EHRFL framework significantly enhances the practical applicability of federated learning in clinical settings, enabling healthcare institutions to build cost-effective, institution-specific predictive models within budget constraints.
Table 1: Comparison with prior EHR federated learning studies. A common limitation of prior work is the oversimplification of federated learning client construction, where a single EHR dataset is artificially partitioned to simulate multiple clients, or multiple hospitals with identical EHR schemas are used, failing to reflect real-world EHR heterogeneity.
Work Objective Dataset(s) Dataset Heterogeneity
[18] In-hospital mortality prediction MIMIC-III No
[19] 7-day mortality prediction for COVID-19 patients Mount Sinai Health System No
[20] In-hospital mortality and AKI prediction in ICU patients eICU No
[21] AKI and sepsis risk prediction in ICU eICU No
Ours Institution-specific clinical prediction modeling MIMIC-III (CV/MV), MIMIC-IV, eICU (South/West) Yes
Table 2: Comparison with prior studies using client similarity measurement in federated learning. Unlike our work, prior studies focus on constructing models optimized for multiple clients simultaneously (e.g., clustered federated learning), rather than building a model targeted for a single institution.
Work Similarity Metric Target Clients
[22] K-Means Clustering Multiple Institutions
[23] Spectral Clustering Multiple Institutions
Ours Cosine Similarity, Euclidean Distance, KL-Divergence Single Institution
@@ 2 Overall Framework
We introduce the two key components of our EHRFL framework. The two components are (1) text-based EHR federated learning (Figure 1) and (2) participating client selection using averaged patient embeddings (Figure 2). The overall UML diagram is shown in Figure 3.
@@
Figure 2: Selection of participating subjects in federated learning based on averaged patient embedding similarity with the host. To ensure privacy, each subject constructs its averaged patient embedding using differential privacy by clipping individual patient embeddings, averaging them, and adding Gaussian noise to the averaged embedding. To ensure consistency in similarity computation, the host applies the same clipping operation to its patient embeddings prior to averaging. Subjects with low similarity scores relative to the host are excluded from the federated learning process.
Figure 3: UML Diagram for EHRFL
@@ 2.1.2 Text-based EHR Linearization
$r_{i}=(e_{i}\oplus n_{i1}\oplus v_{i1}^{\prime}\oplus n_{i2}\oplus v_{i2}^{\prime}\cdots)$ (1)
@@ 2.2 Participating Subject Selection using Averaged Patient Embeddings
In this section, we present our method for selecting suitable participating subjects for the host using averaged patient embeddings. Motivated by prior research [24, 25, 26, 27], we hypothesize that subjects whose data distributions differ significantly from that of the host are more likely to have no positive impact or may even degrade the host model’s performance. To identify suitable subjects, we compute the similarity between the averaged patient embeddings extracted from the text-based EHR data of each candidate subject and that of the host. To ensure patient privacy, we construct each subject’s averaged patient embedding using differential privacy. Specifically, following prior differential privacy implementations, [28, 29, 30], we clip individual patient embeddings, compute their average, and add Gaussian noise to the resulting averaged embedding.
@@ 2.2.1 Subject Selection Process
The subject selection process proceeds as follows. First, the host trains a model on its own data to build a common model for extracting patient embeddings. This model is trained using the text-based EHR modeling method described in Section 2.1. This is to ensure that embeddings are generated within a shared embedding space, despite that host and subjects may have distinct EHR systems. Second, the host shares the trained model weights to each candidate subject. Third, the host and each candidate subject use the model to extract patient embeddings from their respective datasets. Let $l^{h}_{i}\in\mathbb{R}^{d}$ denote the embedding of the $i$-th patient at the host, and $l^{s}_{j}\in\mathbb{R}^{d}$ denote the embedding of the $j$-th patient at a candidate subject. The resulting embeddings for the host are denoted as $\{l^{h}_{1},l^{h}_{2},\dots,l^{h}_{n}\}$, while those for a candidate subject are denoted as $\{l^{s}_{1},l^{s}_{2},\dots,l^{s}_{m}\}$, where $n$ and $m$ represent the number of patients at the host and subject, respectively (note that $l$ is consistent with the notation in Section 2.1.3).
Fourth, both the host and each candidate subject apply $\ell_{2}$-norm clipping to individual patient embeddings to bound the contribution of any single patient. Specifically, each embedding is clipped to a maximum norm $C$ as follows:
$l^{\prime}=l\cdot\min\left(1,\frac{C}{\|l\|_{2}}\right).$ (4)
The resulting clipped patient embeddings for the host are denoted as ($l^{h^{\prime}}_{1}$, $l^{h^{\prime}}_{2}$, …, $l^{h^{\prime}}_{n}$), while the resulting clipped patient embeddings for a single candidate subject are denoted as ($l^{s^{\prime}}_{1}$, $l^{s^{\prime}}_{2}$, …, $l^{s^{\prime}}_{m}$).
Fifth, the host computes its averaged patient embedding from the clipped embeddings as
$\bar{l}^{h^{\prime}}=\frac{1}{n}\sum_{i=1}^{n}l^{h^{\prime}}_{i},$ (5)
and each candidate subject similarly computes an averaged embedding
$\bar{l}^{s^{\prime}}=\frac{1}{m}\sum_{j=1}^{m}l^{s^{\prime}}_{j}.$ (6)
This averaging produces a single embedding vector of dimensionality $d$, where $d$ is the dimensionality of each patient embedding.
Sixth, to ensure patient-level privacy, each candidate subject applies differential privacy to its averaged embedding before sharing it with the host. Following the Gaussian mechanism, random Gaussian noise calibrated to sensitivity $C/m$ is added to $\bar{l}^{s^{\prime}}$, yielding a differentially private averaged embedding:
$\tilde{l}^{s^{\prime}}=\bar{l}^{s^{\prime}}+\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right),$ (7)
where
$\sigma=\frac{C}{m}\cdot\frac{\sqrt{2\log(1.25/\delta)}}{\epsilon},$ (8)
and $(\epsilon,\delta)$ are the differential privacy parameters. In contrast, the host does not add noise to $\bar{l}^{h^{\prime}}$, as the host embedding is not shared externally. However, note that the same clipping operation (in the fourth step) is applied to both host and subject embeddings to ensure consistency in similarity computation.
Seventh, each candidate subject shares their averaged patient embedding, $\tilde{l}^{s^{\prime}}$, to the host. Eighth, for each candidate subject, the host computes the similarity between its own averaged embedding $\bar{l}^{h^{\prime}}$ and the subject’s differentially private averaged embedding. Similarity can be measured using metrics such as cosine similarity, euclidean distance, and KL divergence. Ninth, the host excludes candidate subjects with low similarity scores, proceeding federated learning with subjects of high similarity scores that are not excluded.
@@
A potential concern with this approach is that averaging patient embeddings across clients may result in information loss, which may lead to suboptimal selection of suitable clients. Specifically, given that each client may have many patients, measuring similarity using averaged patient embeddings may lack the necessary granularity. To address this concern, we additionally evaluate subject selection based on patient-wise similarities using non-averaged patient-level embeddings (Supplementary Tables S9, S10). As detailed in Supplementary Tables S7-S10, our experiment results show that the averaged patient embedding method consistently leads to better host performance compared to selecting subjects using non-averaged embedding similarities. Moreover, sharing patient-level embeddings across institutions raises privacy concerns, making the average embedding-based approach a reasonable method from both privacy and performance perspectives.
@@ 2.3 Overall Cost Savings for the Host
[…]
As shown in the total net savings equation, when $K\ll N$, cost savings are substantial since $X$ and $C_{\text{train}}$ scale with (NK)(N-K). Even in the minimal case where K=N1K=N-1, the host still saves approximately $X+C_{\text{train}}$, indicating that our method provides cost benefits even under conservative selection scenarios.
@@
NN Total number of candidate clients in FL (including the host)
@@
$C_{\text{extract}}$ Cost for extracting embeddings from the host’s pre-trained model
$C_{\text{average}}$ Cost for averaging individual embeddings into a single embedding
$C_{\text{embedding}}$ Transmission cost for sending embeddings between servers
$C_{\text{sim}}$ Cost for computing similarity between two embeddings
@@
Extracting embeddings $N\cdot C_{\text{extract}}$
Averaging embeddings $N\cdot C_{\text{average}}$
Transmitting embeddings to the host $N\cdot C_{\text{embedding}}$
@@
Overall Costs $E\cdot C_{\text{train}}+N\cdot(C_{\text{model}}+C_{\text{extract}}+C_{\text{average}}+C_{\text{embedding}}+C_{\text{sim}})$
@@
Total Net Savings $(N-K)\cdot X+((N-K)\cdot R\cdot L-E)\cdot C_{\text{train}}$ $(2\cdot(N-K)\cdot R-N)\cdot C_{\text{model}}$ $N\cdot(C_{\text{extract}}+C_{\text{average}}+C_{\text{embedding}}+C_{\text{sim}})$
@@ 3.1 Datasets
In summary, we construct a total of five clients, representing four distinct EHR database systems, covering three different geographical regions. Detailed demographic information and label distributions for each client are provided in Supplementary Tables S2 and S3.
@@ 3.2.1 Cohort & Prediction Tasks
We evaluate model performance on 12 clinical prediction tasks, following the tasks in [2]. A diverse set of tasks is included to demonstrate the effectiveness of our framework in building host models across various clinical scenarios. The default prediction window is 24 hours, unless otherwise specified (see Supplementary Table S1 for detailed task descriptions). Each model is trained using a multi-task learning, with 12 distinct prediction heads each corresponding to the individual tasks. The data preprocessing code for cohort filtering and prediction task labeling is provided at: https://github.com/Jwoo5/integrated-ehr-pipeline/tree/federated.
3.2.2 Differential Privacy Parameters
Individual patient embeddings are clipped with an $\ell_{2}$-norm threshold of $C=1.0$ prior to averaging. The privacy budget is set to $\epsilon=1.0$ and the failure probability to $\delta=10^{-5}$.
@@ 3.2.4 Training Procedure
Since the objective of our framework is to develop a model tailored specifically for the host, the final host model selected is the one that achieves the highest validation performance for the host across all communication rounds. The experiments were conducted using four NVIDIA RTX 3090 GPUs with CUDA version 11.4, with each host model requiring approximately 1–2 days for training. Detailed reproduction steps along with the full training code are provided in the following repository: https://github.com/ji-youn-kim/EHRFL.
@@ 4 Results
Table 4: Comparison of the host performance (average macro AUROC of the 12 tasks) trained solely on local data (i.e., Single) versus trained on federated learning using text-based EHR modeling with a fixed number of clients. The standard deviation results are listed in Supplementary Table S4. Each client column represents the performance when the client is the host. Average and Best represent the average and best host performance across all subject combinations for the fixed number of clients. Bold indicates federated learning performance higher than Single results.
@@ 4.1 Federated Learning with Text-based EHR Modeling
In this section, we evaluate the effectiveness of federated learning using text-based EHR modeling. Table 4 shows a comparison of the host model performance when trained solely on its own data (i.e., Single) versus when trained using federated learning with text-based EHR modeling. The standard deviation results for Table 4 are listed in Supplementary Table S4. The performance of each host show that federated learning with text-based EHR modeling generally leads to improved host performance compared to training solely on local data (see bolded values). These results indicate that text-based EHR federated learning enables leveraging data from multiple clients to improve the host’s model performance, despite heterogeneity in EHR systems. This trend is consistent across all of the four federated learning algorithms evaluated. Importantly, text-based EHR modeling does not require extensive preprocessing or expert-driven domain-specific mappings, enabling low-cost federated learning when building a model for the host using heterogeneous EHR systems.